OpenStudy (anonymous):

Expand and simplify: [(2x+3)+y][(2x+3)-y]

OpenStudy (anonymous):

\((a+b)(a-b)=a^2-b^2\) with \(a=2x+3,b=y\)

OpenStudy (anonymous):

thanks!